OpenStudy (kaylala):

TOPIC: (trigonometry) Addition and Subtraction Identity FOrmulas Find the value of: sin(-105)

OpenStudy (kaylala):

@agent0smith

OpenStudy (kaylala):

@Loser66

OpenStudy (kaylala):

@Compassionate

OpenStudy (agent0smith):

105 = 60 + 45 and sin(-x) = -sinx so sin(-105) = -sin(60+45)

OpenStudy (kaylala):

is the answer = (2(sqrt6))/4 ? @agent0smith

OpenStudy (agent0smith):

I don't know, wolfram alpha could check for you. Just use the angle sum formula

OpenStudy (kaylala):

i did it but i'm not sure with the signs